OpenStudy (anonymous):

I'm new to this website and how it works, but I would appreciate it if somebody helped me! Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7.

OpenStudy (phantomcrow):

What did you notice when you graphed these two points on a plane?

OpenStudy (anonymous):

@Phantomcrow It opens to the right?

OpenStudy (phantomcrow):

Yes! And you know what the standard equation of a parabola that opens to the right is?

OpenStudy (anonymous):

@Phantomcrow Can you help me out with that one

OpenStudy (anonymous):

It is 3, you're welcome.

OpenStudy (phantomcrow):

Sure.

OpenStudy (phantomcrow):

Since the vertex is at the origin, there is no need to add any constants for shifts. The graph is simply \[y^2=x\]

OpenStudy (anonymous):

@Phantomcrow By the way the Answer choices are y= 1/28x^2 x= 1/28y^2 -28y= x^2 y^2=14x What confuses me is how these tie to the standard formula

OpenStudy (phantomcrow):

D. Despite it having an 'x' term, its vertex is still at the origin.

OpenStudy (phantomcrow):

Standard form is one of the ways to write the graph of a parabola. It is commonly seen as:\[ax^2+bx+c\]

OpenStudy (anonymous):

@Phantomcrow Wow thanks, would it be cool if you helped me with one more?

OpenStudy (phantomcrow):

Sure

OpenStudy (anonymous):

@Phantomcrow Find the standard form of the equation of the parabola with a focus at (0, 8) and a directrix at y = -8, i'm guessing this is similar or even exact to the last one, so it opens to the left

OpenStudy (phantomcrow):

Actually, it does not open to the left this time. You are given y values for focus and directrix, not x values.

OpenStudy (anonymous):

@Phantomcrow So y values open to the right and x values open to the left?

OpenStudy (phantomcrow):

Y values lie across a vertical line, you your parabola will be opening upwards. You can easier see this if you look at the (0.8) and y= -8 on a graph. See that the they are equidistant from a point.

OpenStudy (anonymous):

@PhantomCrow you graph (0,8) correct?

OpenStudy (phantomcrow):

Yes. Graph the point (0,8) and y= -8 and you should be able to visually see the the point that is equidistant from them. There is a formula for finding distance but I believe simply looking at the graph for this problem will provide you with the answer.

OpenStudy (anonymous):

@PhantomCrow The choices are y= 1/32x^2 y^2= 8x y^2=32x y=1/8x^2 Would it be B?

OpenStudy (phantomcrow):

Oh, I see.